# Do MLLMs Really Understand the Charts?

### v1

@@ Abstract
Although Multimodal Large Language Models (MLLMs) have demonstrated increasingly impressive performance in chart understanding, most of them exhibit alarming hallucinations and significant performance degradation when handling non-annotated charts [^1] The non-annotated charts are those that require viewers to estimate values using the vertical/horizontal axis scale.. Therefore, a question arises: Do MLLMs really understand the charts? Since a human is capable of understanding charts and estimating the values by visual reasoning, we first carefully establish a comprehensive Chart Reasoning Benchmark (CRBench) to rigorously evaluate the visual reasoning abilities of MLLMs on non-annotated charts. We argue that MLLMs are primarily relying on recognition rather than reasoning to interpret the charts. To steer MLLMs to reasonable chart understanding, we propose ChartReasoner that mimics human behavior by grounding their estimation in chart understanding. Extensive results on the proposed CRBench show that ChartReasnoner-3B/7B achieves superior performance in chart reasoning, even compared to GPT-4o and Gemini-2.5-Flash. More importantly, ChartReasnoner also demonstrates the visual reasoning abilities in general chart comprehension on public benchmarks, leading to significant performance gains and enabling MLLMs to rationally understand the charts. The code and dataset will be publicly available upon publication.
Introduction
@@
Multimodal Large Language Models (MLLMs) (Bai et al. 2025; Comanici et al. 2025; OpenAI et al. 2024a; Lu et al. 2024) now play a pivotal role in the field of Artificial Intelligence, particularly for understanding complex visual data. These models have demonstrated a remarkable ability to process charts, analyze their content, provide insightful explanations, and achieve competitive performance on many existing chart benchmarks (Wang et al. 2024; Masry et al. 2022; Xu et al. 2024b; Masry et al. 2025a; Xia et al. 2025).
However, we discover an intriguing phenomenon: When specific numerical annotations are missing in the charts, in other words, when handling non-annotated charts, the MLLMs’ propensity for hallucination increases dramatically—-a failure exemplified in Figure 1, where a model defaults to hallucination in non-annotated charts while our proposed method succeeds through multi-step reasoning. This leads us to a fundamental question: Do MLLMs really understand the charts?
@@
Figure 1: illustration of the visual reasoning deficit in MLLMs when processing non-annotated charts. A “standard Initial Model”, limited by its underdeveloped visual reasoning capacity, often resorts to guessing and fails. In contrast, our “Optimized Model” (ChartReasoner) executes a deliberate, human-like reasoning chain—identifying the target, locating data based on visual scales, and forming a conclusion—to successfully estimate the value.
To investigate this issue, we elaborately introduce a Chart Reasoning Benchmark (CRBench), a novel benchmark that specifically requires visual reasoning for chart understanding, focused on the task of numerical estimation on non-annotated charts. Our goal was to rigorously evaluate the visual reasoning capabilities of current MLLMs on chart understanding and explore whether they truly comprehend the visual clues in the charts. The evaluation reveals that not only open-source MLLMs (Bai et al. 2025; Zhu et al. 2025; Lu et al. 2024) face performance degradation, but even powerful close-source MLLMs, such as GPT-4o (OpenAI et al. 2024a) and Gemini-2.5-Flash (Comanici et al. 2025), also struggle significantly with CRBench. Consequently, we argue that most MLLMs are much more relying on visual recognition capabilities, e.g., Optical Character Recognition (OCR), to read the charts, rather than employing genuine visual reasoning to truly understand them.
Inspired by the success of Reinforcement Learning (RL) in enhancing textual reasoning in domains like mathematics and coding (DeepSeek-AI et al. 2025; OpenAI et al. 2024b), we introduce ChartReasoner, a new model forged through a novel Reinforcement Finetuning (RFT) strategy. This two-stage framework is designed not just to recognize the charts, but also understand the charts via visual reasoning. It begins with a “cold start” through Chain-of-Thought Supervised Fine-Tuning (CoT-SFT) to activate a structured thinking paradigm within the MLLM. This initial ability is then dramatically amplified in the second stage, where the Group Relative Policy Optimization (GPRO) (Shao et al. 2024) training method selectively reinforces correct reasoning pathways and penalizes flawed logic. The experimental results demonstrate that this strategy significantly enhances the visual reasoning capabilities of ChartReasoner, achieving superior performance on the CRbench. Besides, we discovered that the foundational visual reasoning skills cultivated by our method generalize broadly, leading to substantial performance gains across a diverse suite of public chart understanding benchmarks. CharReasoner exhibits highly potential in moving towards a more genuine form of chart comprehension.
The main contributions are summarized as follows:
@@
  • We introduce CRBench, a distinctive benchmark designed to isolate and evaluate genuine visual reasoning capability in chart understanding. It overcomes the limitations of prior work by focusing exclusively on numerical estimation tasks, thus disentangling reasoning from text recognition.
@@
  • We propose ChartReasoner, a powerful MLLM with significantly enhanced visual reasoning capabilities for chart understanding. It achieves excellent performance on our challenging CRBench benchmark, compared with chart-specific and general MLLMs, even surpassing powerful proprietary models like GPT-4o and Gemini-2.5-Flash.
@@
  • We demonstrate that the visual reasoning ability cultivated by our method is foundational and highly generalizable. ChartReasoner is not confined to the specific numerical estimation task, but achieves substantial performance gains across a diverse suite of public, multi-task chart understanding benchmarks.
Related Work
Chart Understanding
The advent of MLLMs, encompassing both powerful open-source and close-source (Comanici et al. 2025; OpenAI et al. 2024a; Team et al. 2025), has laid a foundational groundwork for automated chart understanding. Driven by large-scale visual instruction tuning (Liu et al. 2023), this progress has spurred the development of numerous chart-specific fine-tuned models (Han et al. 2023; Zhang et al. 2024; Masry et al. 2024, 2025b) tailored to enhance performance on these specialized tasks. However, the predominant training paradigm for these models relies on SFT (Liu et al. 2023; Wei et al. 2022), which, as we argue, can lead to superficial mimicry rather than genuine reasoning. In contrast, our work explores a paradigm based on RL (Ouyang et al. 2022) to cultivate a deeper, more robust form of chart understanding.
Figure 2: The training paradigm of ChartReasoner and the data distribution of CRBench. A: ChartReasoner leverages a two-stage RFT strategy. Stage 1 activates the model’s reasoning abilities via SFT on CoT data, while Stage 2 uses GRPO with a multi-component reward system to reinforce correct chart understanding. B: The composition of CRBench, detailing the distribution of seven chart types across both synthetic and real data sources.
Methods Synthetic Charts Real Charts Overall
Box Area Radar Scatter Bar Line Combo Bar Line Combo
Human 74.72 56.70 83.97 77.90 90.10 93.71 77.27 80.36 65.22 52.00 78.64
Open-source Models
InternVL3-2B (Zhu et al. 2025) 31.87 7.32 5.77 23.19 51.40 45.70 32.98 39.29 33.04 20.80 31.59
Qwen2.5-VL-3B (Bai et al. 2025) 49.45 15.24 19.23 52.74 74.58 82.45 62.48 67.86 54.78 48.80 57.68
Ovis1.6-llama3.2-3B (Lu et al. 2024) 16.48 2.44 10.90 14.22 17.04 12.25 11.52 12.50 13.04 9.60 12.47
Gemma-3-4B (Team et al. 2025) 15.38 3.05 10.90 10.28 8.38 8.94 6.28 5.36 9.57 8.80 8.32
Qwen2.5-VL-7B (Bai et al. 2025) 72.53 28.05 26.92 66.08 84.08 86.42 71.03 63.39 60.87 47.20 66.25
InternVL3-8B (Zhu et al. 2025) 53.85 19.51 14.10 43.33 62.29 61.92 48.52 58.93 55.65 42.40 42.40
Close-source Models
GPT-4o (OpenAI et al. 2024a) 32.97 15.24 12.82 25.82 22.35 25.17 19.90 15.18 26.96 16.80 21.69
Gemini-2.5-Flash (Comanici et al. 2025) 72.53 32.93 11.54 59.96 70.11 71.19 55.50 53.57 46.96 37.60 55.32
Chart-Specific Models
ChartGemma-3B (Masry et al. 2025b) 10.99 10.98 7.05 21.44 42.74 32.78 24.43 37.50 42.61 28.80 26.74
TinyChart-3B (Zhang et al. 2024) 13.19 7.93 7.69 25.16 56.15 54.30 36.65 57.14 40.87 34.40 35.83
ChartInstruct-7B (Masry et al. 2024)
10.99 1.22 4.49 16.63 35.47 16.56 18.85 51.79 45.22 18.40 20.91
ChartVLM-7.3B (Xia et al. 2025) 9.89 10.37 7.69 10.28 70.39 45.36 32.81 50.00 54.78 35.20 33.63
ChartLlama-13B (Han et al. 2023) 10.99 1.83 5.77 5.25 3.35 4.30 4.01 8.04 9.57 7.20 5.01
Table 1: Performance evaluation on CRBench with direct answering, covering a range of open-source, close-source, and chart-specific models. This table shows the results when models are prompted to directly output the final answer without intermediate reasoning steps.
A suite of chart benchmarks, including ChartQA (Masry et al. 2022), PlotQA (Methani et al. 2020), and others (Xu et al. 2024b; Han et al. 2023; Wang et al. 2024), have been proposed to evaluate MLLMs on diverse chart tasks. While instrumental, their task design often prioritizes language-level reasoning (Lewkowycz et al. 2022) (e.g., calculating averages or summarizing) over fundamental visual reasoning (Johnson et al. 2017; Hudson and Manning 2019). The core skill of grounding reasoning in a chart’s geometry, such as estimating values from axes on non-annotated charts, remains largely untested. Our CRBench is specifically designed to isolate and evaluate this crucial visual reasoning capability.
Visual Reasoning
@@
RL has been successfully employed to significantly enhance the reasoning abilities of Large Language Models (LLMs) (DeepSeek-AI et al. 2025), particularly in complex domains like mathematics and coding (Shao et al. 2024). Inspired by this success in language-only models, several works have begun to apply similar RL-based paradigms to MLLMs (Feng et al. 2025; Tan et al. 2025; Huang et al. 2025), successfully activating their visual reasoning capabilities on various challenging tasks (e.g., visual counting, spatial transformation). Building on these advancements, our work adapts and applies an RL-based framework to the specialized domain of chart understanding.
Chart Reasoning Benchmark
@@
We introduce Chart Reasoning Benchmark (CRBench), a comprehensive, multi-domain, and reasoning-centric benchmark designed to rigorously assess the visual interpretation capabilities of MLLMs on charts that lack explicit numerical annotations. Engineered to move beyond simple OCR-dependent tasks, the benchmark comprises a total of 2,453 question-answer pairs. It features a majority (2,101 pairs) of synthetically generated charts to ensure controlled complexity and a significant portion (352 pairs) sourced from real-world examples to guarantee practical relevance. The benchmark provides extensive coverage across seven primary chart types, including bar, line, scatter, and combo charts, with a detailed breakdown of the data distribution shown in Figure 2. This structural diversity is complemented by thematic breadth, with data spanning 38 distinct topics, including finance, healthcare, and technology, etc. This dual emphasis on structural and thematic variety ensures a rigorous evaluation, mitigating the risk of models overfitting to specific chart formats or familiar domains.
Data Curation
Our data generation process integrates automated synthesis with manual curation to ensure both scale and quality.
@@
Our synthetic chart generation process is partially adapted from the Code-as-Intermediary Translation (CIT) methodology proposed by (He et al. 2024), where executable plotting code serves as the ground truth for each chart. The process begins with a curated set of seed scripts, which are then programmatically diversified using Self-Instruct (Wang et al. 2023) and Evol-Instruct (Xu et al. 2024a) techniques to generate a vast library of visually complex charts. A critical constraint is the deliberate omission of numerical labels on data points, ensuring that every chart necessitates visual estimation. To maximize yield, a self-repair mechanism leverages an LLM to debug and correct any code that fails during execution. Following an automated visual fidelity check by a multimodal language model, the entire collection of synthesized data underwent a final, rigorous human review. This manual verification step served to confirm the high quality of the chart images and the accuracy of their corresponding question-answer pairs. This code-centric approach, combined with multiple stages of validation, provides an unimpeachable ground truth, allowing us to generate verifiably correct Q&A pairs.
@@
Evaluation Protocol
Standard exact-match accuracy is ill-suited for value estimation from non-annotated charts, as it fails to account for the slight perceptual ambiguity inherent in the task, even for human observers. To address this, we employ a relaxed accuracy metric, which judges a prediction correct if its relative error from the ground-truth value falls within a tolerance threshold, denoted as $\tau$. To align this threshold with human performance, we conducted an empirical study and found that human estimations consistently fall within a 2% error margin. Accordingly, we empirically set $\tau=0.02$. Formally, a model’s predicted value, $A_{pred}$, is deemed correct if and only if it satisfies the following condition relative to the ground truth, $A_{gt}$:

### v2

@@ Abstract
Although Multimodal Large Language Models (MLLMs) have demonstrated increasingly impressive performance in chart understanding, most of them exhibit alarming hallucinations and significant performance degradation when handling non-annotated charts [^1] The non-annotated charts are those that require viewers to estimate values using the vertical/horizontal axis scale.. We argue that current MLLMs rely largely on visual recognition rather than visual reasoning to interpret the charts, and visual estimation of numerical values is one of the most fundamental capabilities in chart understanding that require complex visual reasoning. To prove this, we introduce ChartVRBench, a benchmark meticulously designed to isolate and evaluate visual reasoning ability in chart understanding. Furthermore, we propose ChartVR-3B/7B trained with a novel Visual Reasoning Reinforcement Finetuning (VR-RFT) strategy to strengthen genuine chart visual reasoning abilities. Extensive experiments show that ChartVR achieves superior performance on ChartVRBench, outperforming even powerful proprietary models. Moreover, the visual reasoning skills cultivated by the proposed VR-RFT demonstrate strong generalization, leading to significant performance gains across a diverse suite of public chart understanding benchmarks. The code and dataset will be publicly available upon publication.
1 Introduction
@@
Multimodal Large Language Models (MLLMs) (bai2025qwen2; comanici2025gemini; hurst2024gpt; lu2024ovis) now play a pivotal role in the field of Artificial Intelligence, particularly for understanding complex visual data. These models have demonstrated a remarkable ability to process charts, analyze their content, provide insightful explanations, and achieve competitive performance against existing chart benchmarks (wang2024charxiv; masry2022chartqa; xu2023chartbench; masry2025chartqapro; xia2024chartx).
Estimating numerical values from charts is a fundamental capability in chart understanding that involves interpreting visual representations to extract or approximate the underlying numbers. The core principle is to understand the mapping between the visual elements (e.g., the position, length, or angle of a mark) on the chart and the data scale it represents. However, when specific numerical annotations are missing from the chart, the propensity of MLLMs to hallucination increases dramatically (xu2023chartbench), as exemplified in Figure 1. This leads us to a fundamental question: Do MLLMs really understand the charts?
This failure suggests that current MLLMs excel at recognizing about textual content within charts but struggle profoundly with reasoning from their underlying visual geometry. We argue that it stems from the fundamental reliance of MLLMs on textual recognition over genuine visual reasoning. To systematically diagnose this core ability, we introduce the Chart Visual Reasoning Benchmark (ChartVRBench), which is meticulously designed to isolate numerical value estimation on non-annotated charts, forcing models to move beyond textual recognition. The evaluation reveals that not only open-source MLLMs (bai2025qwen2; zhu2025internvl3; lu2024ovis) face performance degradation, but even powerful close-source MLLMs, such as GPT-4o (hurst2024gpt) and Gemini-2.5-Flash (comanici2025gemini), also struggle significantly with ChartVRBench.
@@
Figure 1: Illustration of the visual reasoning deficit in MLLMs when processing non-annotated charts. A standard model, limited by its underdeveloped visual reasoning capacity, often resorts to guessing and fails. In contrast, our ChartVR executes a deliberate, human-like reasoning chain—identifying the target, locating data based on visual scales, and forming a conclusion—to successfully estimate the value.
Moreover, inspired by the success of Reinforcement Learning (RL) in enhancing textual reasoning for mathematics and coding (guo2025deepseek; jaech2024openai; tan2025reason; huang2025vision), we propose ChartVR, a series of MLLMs forged with a novel Visual Reasoning Reinforcement Finetuning (VR-RFT) strategy to strengthen genuine chart visual reasoning abilities. The first stage, Visual Reasoning Activation, uses a Chain-of-Thought Supervised Finetuning (CoT-SFT) (liu2023visual) to compel the model to externalize a step-by-step analysis of the chart’s visual components. This forces the model to learn an explicit protocol for geometric interpretation, such as locating axes and grounding queries to graphical marks, thereby forming the structural foundation of its visual reasoning capability. Building on this, the second stage, Visual Reasoning Generalization, employs Group Relative Policy Optimization (GRPO) (shao2024deepseekmath) to further refine this process. By training on a curated dataset of ambiguous samples where the initial model’s judgment is inconsistent, we force it to make finer perceptual discriminations. This training process is guided by a novel continuous accuracy reward function with a quadratic formulation, providing a dense signal directly proportional to the accuracy of the visual estimation. In summary, these stages steer ChartVR to a robust, generalizable visual reasoning capability for charts.
The extensive experiments demonstrate that ChartVR achieves superior performance on ChartVRBench, even comparable to powerful proprietary models like Gemini-2.5-Flash (comanici2025gemini). More importantly, we demonstrate that the foundational skill cultivated by our method is highly generalizable. ChartVR exhibits significant performance gains across a diverse suite of public, multi-task chart understanding benchmarks (wang2024charxiv; xu2023chartbench; masry2025chartqapro), proving the effectiveness of our approach in building more rational and reliable MLLMs for chart comprehension.
The main contributions of this work are summarized as follows:
@@
  • We introduce ChartVRBench, a distinctive benchmark designed to isolate and evaluate genuine visual reasoning capability in chart understanding. It overcomes the limitations of prior work by focusing exclusively on numerical estimation tasks, thus disentangling reasoning from text recognition.
@@
  • We propose ChartVR, a series of MLLMs with significantly enhanced visual reasoning capabilities for chart understanding. It achieves excellent performance on our challenging ChartVRBench, compared with chart-specific and general MLLMs, even surpassing powerful proprietary models like Gemini-2.5-Flash.
@@
  • We demonstrate that the visual reasoning ability cultivated by our method is foundational and highly generalizable. ChartVR is not confined to the specific numerical estimation task, but achieves substantial performance gains across a diverse suite of public, multi-task chart understanding benchmarks.
2 Related Work
2.1 Chart Understanding Benchmarks
A suite of benchmarks has been developed to evaluate the chart comprehension capabilities of MLLMs. Early benchmarks, such as ChartQA (masry2022chartqa) and PlotQA (methani2020plotqa), primarily focused on descriptive tasks. More recently, benchmarks like CharXiv (wang2024charxiv), ChartQAPro (masry2025chartqapro), and ChartMuseum (tang2025chartmuseum) have raised the bar by incorporating complex questions and diverse, real-world charts. While these works encompass a wide range of tasks, they often conflate general reasoning with the core challenge of visual interpretation. The most related work to ours is ChartBench (xu2023chartbench); while it also focuses on non-annotated charts, it is composed of mostly synthetic data with limited visual diversity. Similarly, recent work by mukhopadhyay2024unraveling revealed critical flaws in the consistency and robustness of MLLMs but stopped short of attributing these shortcomings to a fundamental deficit in visual reasoning. We argue this deficit—the core skill of visual reasoning in a chart’s geometry, such as numerical value estimation—remains largely untested. Our ChartVRBench is specifically designed to isolate and evaluate this crucial visual reasoning capability.
2.2 Chart Understanding with MLLMs
Many general-purpose MLLMs, such as gpt-4o (hurst2024gpt), Gemini-2.5 Series (comanici2025gemini), and Qwen (bai2025qwen2), are increasingly applied to chart understanding tasks. In parallel, the development of specialized Chart MLLMs has been rapid, with many models like ChartLlama (han2023chartllama) and ChartGemma (masry2024chartgemma). However, their development has largely depended on SFT, a paradigm that, as we argue, tends to cultivate superficial recognition at the expense of genuine reasoning. Recognizing this, a recent wave of models (chen2025chartr1; masry2025bigcharts), have incorporated RL to enhance complex, multi-step reasoning. While these RL-based approaches represent a significant step forward, their training objectives often prioritize the final accuracy of text-heavy queries, which can leave the foundational skill of visual grounding underdeveloped. In contrast, our ChartVR is specifically designed to address this fundamental layer. Its RFT framework is meticulously crafted to cultivate the core ability to reason directly from visual geometry, aiming to develop a genuine visual reasoning capability rather than optimizing the textual reasoning that typically follows.
2.3 Reasoning in Chart Understanding
@@
Reinforcement Learning (RL) has been successfully employed to enhance the reasoning abilities of Large Language Models (LLMs), allowing them to move beyond the static data distributions of SFT (ouyang2022training). By learning from reward feedback, models have shown significant improvements in complex domains like mathematics and coding (guo2025deepseek; shao2024deepseekmath). Inspired by this success, several works have begun to apply similar RL-based paradigms to MLLMs (feng2025video; tan2025reason; huang2025vision), activating their visual reasoning on tasks like visual counting and spatial transformation. Building on these advancements, our work adapts this powerful paradigm to the specialized domain of chart understanding.
Figure 2: The training paradigm of ChartVR and the data distribution of ChartVRBench. A: ChartVR leverages a two-stage RFT strategy. Stage 1 activates the model’s reasoning abilities via SFT on CoT data, while Stage 2 uses GRPO with a multi-component reward system to reinforce correct chart understanding. B: The composition of ChartVRBench, detailing the distribution of seven chart types across both synthetic and real data sources.
3 ChartVRBench
@@
We introduce Chart Visual Reasoning Benchmark (ChartVRBench), a comprehensive, multi-domain, and reasoning-centric benchmark designed to rigorously assess the visual interpretation capabilities of MLLMs on charts that lack explicit numerical annotations. Engineered to move beyond simple OCR-dependent tasks, the benchmark comprises a total of 2,453 question-answer pairs. It features a majority (2,101 pairs) of synthetically generated charts to ensure controlled complexity and a significant portion (352 pairs) sourced from real-world examples to guarantee practical relevance.
The benchmark provides extensive coverage across seven primary chart types, including bar, line, scatter, and combo charts, with a detailed breakdown of the data distribution shown in Figure 2. This structural diversity is complemented by thematic breadth, with data spanning 38 distinct topics, including finance, healthcare, and technology. This dual emphasis on structural and thematic variety ensures a rigorous evaluation, mitigating the risk of models overfitting to specific chart formats or familiar domains.
While existing benchmarks have significantly advanced the field, they predominantly focus on general high-level Question Answering (QA), where visual reasoning is often conflated with textual extraction (OCR) and logical reasoning. ChartVRBench fills a critical gap by strictly isolating the visual reasoning capability of numerical value estimation on non-annotated charts, preventing models from relying on text recognition shortcuts.To clearly demonstrate how our benchmark compares to contemporary works, we present a feature-wise comparison in Table 1.
Table 1: Comparison between ChartVRBench and existing representative chart QA benchmarks. Symbols: ✓ Fully Supported / High Quality; $\triangle$ Partially Supported / Mixed; $\times$ Not Supported / Low Quality.
Feature
EvoChart
huang2025evochart
ChartBench
xu2023chartbench
CharXiv
wang2024charxiv
ChartQAPro
masry2025chartqapro
ChartMuseum
tang2025chartmuseum
Ours
Real-World Charts $\times$
Broad Topic Coverage $\times$ $\times$ $\triangle$ $\triangle$ $\triangle$
Non-Annotated Charts $\times$ $\times$ $\times$ $\times$
Isolates Visual Reasoning $\times$ $\triangle$ $\times$ $\times$
3.1 Data Curation
@@
Our synthetic chart generation process is partially adapted from the Code-as-Intermediary Translation (CIT) methodology proposed by he2024distill, where executable plotting code serves as the ground truth for each chart. The process begins with a curated set of seed scripts, which are then programmatically diversified using Self-Instruct (wang2022self) and Evol-Instruct (xu2024wizardlm) techniques to generate a vast library of visually complex charts. A critical constraint is the deliberate omission of numerical labels on data points, ensuring that every chart necessitates visual estimation. To maximize yield, a self-repair mechanism leverages an LLM to debug and correct any code that fails during execution. Following an automated visual fidelity check by a MLLM, the entire collection of synthesized data underwent a final, rigorous human review. This manual verification step served to confirm the high quality of the chart images and the accuracy of their corresponding question-answer pairs. This code-centric approach, combined with multiple stages of validation, provides an unimpeachable ground truth, allowing us to generate verifiably correct Q&A pairs.
@@
3.2 Evaluation Protocol
Standard exact-match accuracy is ill-suited for value estimation from non-annotated charts, as it fails to account for the slight perceptual ambiguity inherent in the task, even for human observers. To address this, we employ a relaxed accuracy metric, which judges a prediction correct if its relative error from the ground-truth value falls within a tolerance threshold, denoted as $\tau$. To align this threshold with human performance, we conducted an empirical study and found that human estimations consistently fall within a 2% error margin. Accordingly, we empirically set $\tau=0.02$.
Formally, a model’s predicted value, $A_{pred}$, is deemed correct if and only if it satisfies the following condition relative to the ground truth, $A_{gt}$: